# Strange stars admixed with dark matter: equiparticle model in a two fluid approach

### v2

@@ V Summary and concluding remarks
In this work we have investigated the possible scenarios concerning the existence of strange quark stars with dark matter in its composition. The description of ordinary strongly interacting matter was made through the so called equiparticle (EQP) model [33, 35], in which the constituent quark masses ($u$, $d$, $s$ quarks) are affected by the medium by means of suitable baryonic density dependent functions. The parameters present in these functions were chosen from the analysis of the stellar matter stability window, by imposing parametrizations with lower energy per baryon in comparison with the iron nuclei. Similar studies were performed in [87, 97], but based on versions of the MIT bag model in the description of the “visible” sector of the system. An improved version of the EQP model was proposed in [52, 34], where the Polyakov loop was included with the aim of representing the gluonic dynamics of the strong interaction. As a consequence, the new model was verified to be able of describing quark matter (first order) transition from confined to deconfined phase at zero-temperature regime.
@@
The dark matter part of the system was studied from two different models, namely, a fermionic and a bosonic one. The former was described by a Lagrangian density in which the dark spinor feels a repulsive interaction coming from a dark vector field. The full equations of state only depend on two free parameters: the dark fermion mass ($m_{\chi}$), and the ratio between the strength of the vector interaction ($g_{V}$) and the vector field mass ($m_{V}$). We consider these quantities inside the ranges of $0.5~{}\mbox{GeV}\leqslant m_{\chi}\leqslant 4.5~{}\mbox{GeV}$ [84] and $0.1~{}\mbox{fm}\leqslant C_{V}=g_{V}/m_{V}\leqslant 5~{}\mbox{fm}$ [71]. With regard to the bosonic model, we consider a version where we neglect interactions between the dark boson and the quarks. As in the fermionic case, we also take into account a vector interaction, in this case crucial in order to avoid the collapse of the resulting star due to the absence of degeneracy pressure. The final equations for energy density and pressure are derived and, once again, two free parameters are needed to completely determine the model: the dark scalar mass ($m_{\sigma}$) and $C_{\sigma\phi}=g_{\sigma}/m_{\sigma}$, where $g_{\sigma}$ is the interaction strength of the dark scalar complex field with the dark vector field. Here we take values compatible with $10^{-2}~{}\mbox{MeV}\leqslant m_{\sigma}\leqslant 10^{8}~{}\mbox{MeV}$ and $10^{-2}~{}\mbox{MeV}^{-1}\leqslant C_{\sigma\phi}\leqslant 10^{3}~{}\mbox{MeV}^{-1}$ [79].
@@
Since luminous and dark matter models are defined, we let both sectors interact only by gravity through a two-fluid approach. Our results point out to strange stars admixed with dark matter, with fixed dark matter fraction, as a possible scenario in agreement with astrophysical data coming from the massive millisecond pulsars PSR J0030+0451 [11, 12] and PSR J0740+6620 [13, 14], the NASA’s Neutron star Interior Composition Explorer (NICER) x-ray telescope, and the compact star named as HESS J1731-347 [16].
@@ Acknowledgements
This work is a part of the project INCT-FNA proc. No. 464898/2014-5. It is also supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Grants No. 307255/2023-9 (O. L.), 401565/2023-8 (Universal, O. L.), and doctorade scolarships No. 400879/2019-0 (I. M), and 140324/2024-0 (A. F. C). E. H. R. was supported by the doctorate scholarship from Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES).

### v3

@@ V Summary and concluding remarks
In this work, we have investigated possible scenarios regarding the existence of strange quark stars with dark matter in their composition. The description of ordinary strongly interacting matter was carried out using the so-called equiparticle (EQP) model [40, 42], in which the constituent quark masses ($u$, $d$, $s$ quarks) are influenced by the medium through appropriate baryonic density-dependent functions. The parameters of these functions were chosen based on the analysis of the stellar matter stability window, ensuring parametrizations with lower energy per baryon compared to iron nuclei. Similar studies were conducted in [104, 113], but they relied on versions of the MIT bag model to describe the “visible” sector of the system. An improved version of the EQP model was proposed in [59, 41], incorporating the Polyakov loop to account for the gluonic dynamics of strong interactions. As a result, the new model was found to be capable of describing the first-order phase transition of quark matter from a confined to a deconfined phase in the zero-temperature regime.
@@
The dark matter component of the system was studied from two different models, namely, a fermionic and a bosonic one. The former is described by a Lagrangian density in which the dark spinor experiences a repulsive interaction mediated by a dark vector field. The full equations of state only depend on two free parameters: the dark fermion mass ($m_{\chi}$), and the ratio between the strength of the vector interaction ($g_{V}$) and the vector field mass ($m_{V}$). We consider these quantities within the ranges $0.5~{}\mbox{GeV}\leqslant m_{\chi}\leqslant 4.5~{}\mbox{GeV}$ [80] and $0.1~{}\mbox{fm}\leqslant C_{V}=g_{V}/m_{V}\leqslant 5~{}\mbox{fm}$ [80]. With regard to the bosonic model, we consider a version where we neglect interactions between the dark boson and the quarks. As in the fermionic case, we also take into account a vector interaction, in this case crucial in order to avoid the collapse of the resulting star due to the absence of degeneracy pressure. The final equations for energy density and pressure are derived and, once again, two free parameters are needed to completely determine the model: the dark scalar mass ($m_{\sigma}$) and $C_{\sigma\phi}=g_{\sigma}/m_{\sigma}$, where $g_{\sigma}$ is the interaction strength of the dark scalar complex field with the dark vector field. Here, we adopt values consistent with $10^{-2}~{}\mbox{MeV}\leqslant m_{\sigma}\leqslant 10^{8}~{}\mbox{MeV}$ and $10^{-2}~{}\mbox{MeV}^{-1}\leqslant C_{\sigma\phi}\leqslant 10^{3}~{}\mbox{MeV}^{-1}$ [86].
@@
Since both luminous and dark matter models are defined, we allow the two sectors to interact solely through gravity, following a two-fluid approach. Our results indicate that strange stars admixed with dark matter, with a fixed dark matter fraction, constitute a possible scenario consistent with astrophysical data from massive millisecond pulsars, such as PSR J0030+0451 [16, 15] and PSR J0740+6620 [18, 19], as well as observations from NASA’s Neutron Star Interior Composition Explorer (NICER) x-ray telescope and the compact star HESS J1731-347 [23].
The fourth and fifth observation runs (O4 and O5) of the LIGO-Virgo-KAGRA Collaboration are expected to significantly enhance our understanding of compact stars (CS) by substantially increasing the number of detectable gravitational wave (GW) events, with the potential detection of electromagnetic counterparts [125]. These observations will offer unprecedented opportunities to refine the mass-radius relationship of such objects, providing critical insights into the equation of state of ultra-dense matter. Moreover, the presence of dark matter within these systems could affect GW signals, particularly by modifying post-merger waveform modes [126, 127]. The next generation of ground-based interferometers, such as the Einstein Telescope, will further improve the precision of these measurements, enabling the identification of potential dark matter signatures in binary mergers [128].
Complementing the analysis of these GW events, observational programs such as ATHENA [129], STROBE-X [130], and the Enhanced X-ray Timing and Polarimetry Mission (eXTP) [131] will play an important role in advancing our understanding of CS and their potential DM content. With its high effective area and microsecond timing resolution [132], eXTP, for instance, will enable precise measurements of pulsations and quasiperiodic oscillations providing critical data to determine the mass-radius relationship. Furthermore, eXTP sensitivity to CS surface temperatures will allow for rigorous tests of theoretical predictions regarding DM accumulation inside CS. It is expected that the combination of these data will provide us with enough elements to determine tighter constraints on the mass-radius relationship of CS and probe the presence and the amount of DM accumulated inside these compact objects.
@@ Acknowledgements
This work is part of the project INCT-FNA proc. No. 464898/2014-5. It is also supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Grants No. 307255/2023-9 (O. L.), 401565/2023-8 (Universal, O. L.), and doctorate scholarships No. 400879/2019-0 (I. M), and 140324/2024-0 (A. F. C). E. H. R. was supported by the doctorate scholarship from Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES). O. L. acknowledges M. Dutra for stimulating discussions and valuable comments.